# Optimal Sparsity of Mixture-of-Experts Language Models for Reasoning Tasks

### v2

@@ 1 Introduction
Yet the coefficients of the scaling laws are not universal. Highly expressive models trained under different optimizers or architectures often follow the same loss trajectory but diverge substantially on downstream reasoning benchmarks (Liu et al., 2023). Brandfonbrener et al. (2025) extend the classic laws with loss-to-loss prediction, showing that the mapping between training and test distributions admits its own power law when the distributions differ substantially. These observations imply that optimal budgets must be re-estimated whenever we modify the model or the data pipeline.
@@ 2.2 Scaling Laws of LLMs
Adding MoE experts tends to improve memorization skills more than reasoning skills, motivating new, generalized scaling frameworks that address scaling laws for reasoning performance (Jelassi et al., 2025). scaling laws differ across tasks, the optimal scaling strategy may also vary; for example, knowledge-based QA tasks are “capacity-hungry,” benefiting more from larger model sizes, whereas code-related tasks are “data-hungry,” benefiting more from increased training data (Roberts et al., 2025). Complementing these observations, recent analyses find that models with identical training loss can still exhibit markedly different reasoning performance (GLM-4.5 Team, 2025), highlighting that reasoning skill depends not only on loss but also on how compute and data are allocated.
@@ 3.1 Experimental Setup
Figure 1: Although training and validation loss decrease as the total number of parameters grows, the task loss on GSM8K can sometimes worsen with larger models. Training and validation losses steadily decrease as total or active parameters increase. The HellaSwag task loss follows this scaling trend, whereas GSM8K task loss worsens once total parameters exceed a threshold.
@@ 3.3 Optimal Sparsity for Iso-FLOP Budgets
In Figure 5, we show the task-specific optimal sparsity (i.e. 1-TopK/Experts) against model performance under a fixed FLOPs budget. For memorization benchmarks, lower density (higher sparsity) consistently yields lower task loss and higher accuracy. This pattern aligns with prior studies showing that, when FLOPs are fixed to be constant, sparse models outperform denser models on QA tasks (Abnar et al., 2025). By contrast, on mathematical-reasoning benchmarks, denser models outperform their sparser counterparts. At lower FLOPs, increasing sparsity still reduces loss and improves accuracy; however, once the FLOPs budget grows, denser models begin to perform better, achieving both lower loss and higher accuracy. This shift indicates that reasoning skills admit a compute-dependent optimum, rather than monotonically favoring either sparsity or density.
Figure 5: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, GSM8K and GSM-Plus shift their optima back toward dense models. Task loss (top row) and Accuracy (bottom row) against MoE Density k/E for a fixed active parameter budget.In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently lowers task loss and raises accuracy across all active parameter budgets, in contrast, in the right two tasks (GSM8K, GSM-Plus), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
@@ 3.5 Impact of TTC and Post-Training on Downstream Performance
Test-Time Compute and RL post-training are standard for boosting reasoning on tasks such as mathematical problem solving. We therefore investigated whether the performance trade-offs observed above persist or shift when applying (a) Test-Time Compute (TTC) and (b) RL post-training (GRPO). In Test-Time Compute, we evaluated GSM8K(Cobbe et al., 2021) in a zero-shot setting using Self-Consistency (SC) decoding(Wang et al., 2023), generating $2^{7}$ independent continuations per problem and selecting the most frequent answer. In Post-Training, we fine-tuned each model on the GSM8K training dataset using the GRPO algorithm (Shao et al., 2024). We followed the settings of Zhao et al. (2025) including reward function and fixed the learning rate constant across all model configurations.
@@ 3.6 Coding Task Ablations
Figure 8 summarizes performance as a function of MoE density $k/E$ under matched active compute budgets. As active parameters grow, both HumanEval and MBPP exhibit a clear shift in their optima toward denser configurations: beyond a task-dependent threshold, further increasing sparsity degrades pass@1 despite continued improvements in pre-training loss. This echoes our math findings: when compute allows large active capacity, denser MoE layers yield better procedural reasoning for code synthesis, whereas sparser layers are more favorable only in the low-compute regime. Detailed results for the coding tasks are provided in Appendix C.5.
Figure 8: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, HumanEval and MBPP shift their optima back toward dense models. Accuracy against MoE Density k/E for a fixed active parameter budget.In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently raises accuracy across all active parameter budgets, in contrast, in the right two tasks (HumanEval, MBPP), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
@@ 4 Discussion and Limitations
Our study does not attempt to exhaustively explore all architectural and training settings. The design space of MoE models is vast and it is infeasible to cover every possible combination. Instead, we focused on a representative but systematic sweep over width, expert count, and top-$k$ routing, which already reveals new regularities that prior work did not capture. In particular, we identify an inverted-U relationship between sparsity and reasoning performance that contradicts the monotonic trends often assumed in scaling analyses. Our analysis further shows that this effect can be parsimoniously explained in terms of TPP and active FLOPs: reasoning performance peaks near a task-specific TPP optimum and degrades when models are either too parameter-heavy or too data-starved, even as training loss continues to fall. Existing scaling laws are incomplete and require refinement.
@@ A.2 Post-Training Details
We use GRPO(Shao et al., 2024) with a batch size of 1024, train for 15 epochs, and truncate prompts and generated sequences to 512 and 1024 tokens respectively. The actor’s learning rate is fixed at $5\times 10^{-6}$; the temperature is set to 1.0, the KL-penalty coefficient to $10^{-3}$, and 5 samples are used per prompt. Optimisation employs Adam with $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$, and weight decay of $10^{-2}$. Following Zhao et al. (2025), we implemented a code-execution-based evaluator supporting TinyGSM-style and OpenMathInstruct-1 outputs. For a width of $2048$ with 16 or 64 experts, we swept the learning rate (Fig. 9) and subsequently fixed it to $5\times 10^{-6}$ for all GRPO experiments.
@@ Appendix B Evaluation Setup
We assess logical reasoning capabilities using Mathematical problem-solving is evaluated using GSM8K (Cobbe et al., 2021) with 4-shot prompting and GSM-Plus (Li et al., 2024) with 5-shot CoT prompting. We evaluate comprehension abilities using TriviaQA (Joshi et al., 2017) with 4-shot prompting. Common sense reasoning is assessed through HellaSwag (Zellers et al., 2019) using 4-shot prompting setups. Finally, code reasoning capabilities are benchmarked on HumanEval (Chen et al., 2021) with 0-shot prompting and MBPP (Austin et al., 2021) with 3-shot prompting, both evaluated using the Pass@1 metric. For Test-Time Compute (TTC) experiments specifically, GSM8K evaluation is conducted under a zero-shot setting. To accommodate the variety of valid answer formats, we extend the strict match patterns provided by the lm-evaluation-harness beyond the standard implementation. Our matching criteria accept both the standard GSM8K format (####) and GSM8K-CoT formats prefixed with “The answer is” or “Answer:”.
@@
Dataset TriviaQA HellaSwag GSM8K GSM-Plus HumanEval MBPP
@@
Language EN EN EN EN EN EN
# Instances 17,944 10,042 1,319 10,552 164 378
Few-shot # 4 4 4 (0 for TTC) 5 0 3
Metric Accuracy Accuracy Accuracy CoT Acc. Pass@1 Pass@1
@@ C.3 GSM8K Overfitting Analysis
Figure 15: Performance versus MoE density after removing GSM8K-related training data. Task loss (top) and accuracy (bottom) are plotted against MoE density (k/E) for a fixed active parameter budget. While performance on memorization tasks (TriviaQA, HellaSwag) improves with sparsity, the trend reverses for math reasoning tasks (GSM8K, GSM-Plus) at larger active parameter counts. Dashed segments mark the inverse-scaling regime.
@@ C.5 Detailed Results for Coding Tasks
This appendix provides detailed results for the coding task ablations, mirroring the analyses for mathematical reasoning presented in the main text. Specifically, we detail the non-monotonic relationship between scaling and downstream performance, showing how both task loss (Figure 18, 19) and accuracy (Figure 20) can degrade as pre-training loss improves. We then analyze key architectural factors, including the impact of MoE sparsity (Figure 21) and the optimal Tokens-per-Parameter (TPP) ratio (Figure 22).
@@
Figure 21: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, HumanEval and MBPP shift their optima back toward dense models. Task loss against MoE Density k/E for a fixed active parameter budget.In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently lowers task loss across all active parameter budgets, in contrast, in the right two tasks (HumanEval, MBPP), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.

### v3

# arXiv:2508.18672v3[cs.LG] 01 Mar 2026

@@ 1 Introduction
Yet the coefficients of the scaling laws are not universal. Highly expressive models trained under different optimizers or architectures often follow the same loss trajectory but diverge substantially on downstream reasoning benchmarks (Liu et al., 2023a). Brandfonbrener et al. (2025) extend the classic laws with loss-to-loss prediction, showing that the mapping between training and test distributions admits its own power law when the distributions differ substantially. These observations imply that optimal budgets must be re-estimated whenever we modify the model or the data pipeline.
@@ 2.2 Scaling Laws of LLMs
Adding MoE experts tends to improve memorization skills more than reasoning skills, motivating new, generalized scaling frameworks that address scaling laws for reasoning performance (Jelassi et al., 2025). They provide a theoretical explanation for this asymmetry using graph neural networks. Recent work shows that larger top-$k$ can improve compositional generalization in MoE models (Zhao et al., 2025a), though such studies do not observe the non-monotonic effect of sparsity on reasoning performance that we identify. Since scaling laws differ across tasks, the optimal scaling strategy may also vary; for example, knowledge-based QA tasks are “capacity-hungry,” benefiting more from larger model sizes, whereas code-related tasks are “data-hungry,” benefiting more from increased training data (Roberts et al., 2025). Complementing these observations, recent analyses find that models with identical training loss can still exhibit markedly different reasoning performance (GLM-4.5 Team, 2025), highlighting that reasoning skill depends not only on loss but also on how compute and data are allocated.
@@ 3.1 Experimental Setup
Figure 1: Although training and validation loss decrease as the total number of parameters grows, the task loss on GSM8K can sometimes worsen with larger models. Training and validation losses steadily decrease as total or active parameters increase. The HellaSwag task loss follows this scaling trend, whereas GSM8K task loss worsens once total parameters exceed a threshold. Within each fixed top-k group, moving right on the x-axis corresponds to increasing sparsity (because total experts $E$ increases while $k$ remains fixed), so the right-hand task-loss panels implicitly reflect the same sparsity ordering shown explicitly in Figure 5.
@@ 3.3 Optimal Sparsity for Iso-FLOP Budgets
In Figure 5, we show the task-specific optimal sparsity (i.e. 1-TopK/Experts) against model performance under a fixed FLOPs budget. Because the dataset size is fixed in our experiments, increasing the number of active parameters directly increases the training FLOPs. Thus, as the active parameter count grows, the plots implicitly trace how model performance changes with increasing FLOPs. For memorization benchmarks, lower density (higher sparsity) consistently yields lower task loss and higher accuracy. This pattern aligns with prior studies showing that, when FLOPs are fixed to be constant, sparse models outperform denser models on QA tasks (Abnar et al., 2025). By contrast, on mathematical-reasoning benchmarks, denser models outperform their sparser counterparts. At lower FLOPs, increasing sparsity still reduces loss and improves accuracy; however, once the FLOPs budget grows, denser models begin to perform better, achieving both lower loss and higher accuracy. This shift indicates that reasoning skills admit a compute-dependent optimum, rather than monotonically favoring either sparsity or density.
Figure 5: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, GSM8K and GSM-Plus shift their optima back toward dense models. Task loss (top row) and Accuracy (bottom row) against the ratio of active experts $k$ to total experts $E$ for a fixed active parameter budget. In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently lowers task loss and raises accuracy across all active parameter budgets, in contrast, in the right two tasks (GSM8K, GSM-Plus), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
@@ 3.5 Impact of TTC and Post-Training on Downstream Performance
Test-Time Compute and RL post-training are standard for boosting reasoning on tasks such as mathematical problem solving. We therefore investigated whether the performance trade-offs observed above persist or shift when applying (a) Test-Time Compute (TTC) and (b) RL post-training (GRPO). In Test-Time Compute, we evaluated GSM8K(Cobbe et al., 2021) in a zero-shot setting using Self-Consistency (SC) decoding(Wang et al., 2023), generating $2^{7}$ independent continuations per problem and selecting the most frequent answer. In Post-Training, we fine-tuned each model on the GSM8K training dataset using the GRPO algorithm (Shao et al., 2024). We followed the settings of Zhao et al. (2025b) including reward function and fixed the learning rate constant across all model configurations. Further details regarding the training setup and hyperparameters are provided in Appendix A.2.
@@ 3.6 Coding Task Ablations
Figure 8 summarizes performance as a function of MoE density under matched active compute budgets. As active parameters grow, both HumanEval and MBPP exhibit a clear shift in their optima toward denser configurations: beyond a task-dependent threshold, further increasing sparsity degrades pass@1 despite continued improvements in pre-training loss. This echoes our math findings: when compute allows large active capacity, denser MoE layers yield better procedural reasoning for code synthesis, whereas sparser layers are more favorable only in the low-compute regime. Detailed results for the coding tasks are provided in Appendix C.5.
Figure 8: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, HumanEval and MBPP shift their optima back toward dense models. Accuracy against the ratio of active experts $k$ to total experts $E$ for a fixed active parameter budget. In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently raises accuracy across all active parameter budgets, in contrast, in the right two tasks (HumanEval, MBPP), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
@@ 4 Discussion and Limitations
[…]
@@ 5 Conclusion
6 Reproducibility statement
Due to the anonymity requirements and the design of OpenReview, at the time of ICLR submission, the following resources are made available to the reviewers. All source codes, including those for MoE training and evaluation are provided in the supplementary material. The training data used in this study is publicly available. The model checkpoints are not shared due to their large file sizes, which makes anonymous sharing infeasible. Similarly, while we plan to release the training logs via wandb, maintaining anonymity remains a challenge, so they are not included at this stage.
The training setup is described in Section 3.1 and Appendix A, and the evaluation methodology is provided in Appendix B.
Acknowledgements
[…]
@@ A.2 Post-Training Details
We use GRPO(Shao et al., 2024) with a batch size of 1024, train for 15 epochs totaling 105 steps, and truncate prompts and generated sequences to 512 and 1024 tokens respectively. The actor’s learning rate is fixed at $5\times 10^{-6}$; the temperature is set to 1.0, the KL-penalty coefficient to $10^{-3}$, and 5 samples are used per prompt. Optimisation employs Adam with $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$, and weight decay of $10^{-2}$. Following Zhao et al. (2025b), we implemented a code-execution-based evaluator supporting TinyGSM-style and OpenMathInstruct-1 outputs. For a width of $2048$ with 16 or 64 experts, we swept the learning rate (Fig. 9) and subsequently fixed it to $5\times 10^{-6}$ for all GRPO experiments.
@@ Appendix B Evaluation Setup
We assess logical reasoning capabilities using Mathematical problem-solving is evaluated using GSM8K (Cobbe et al., 2021) with 4-shot prompting and GSM-Plus (Li et al., 2024) with 5-shot CoT prompting. We evaluate comprehension abilities using TriviaQA (Joshi et al., 2017) with 4-shot prompting. Common sense reasoning is assessed through HellaSwag (Zellers et al., 2019) using 4-shot prompting setups. Finally, code reasoning capabilities are benchmarked on HumanEval (Chen et al., 2021) and HumanEval+ (Liu et al., 2023b) with 0-shot prompting and MBPP (Austin et al., 2021) and MBPP+ (Liu et al., 2023b) with 3-shot prompting, both evaluated using the Pass@1 metric. For Test-Time Compute (TTC) experiments specifically, GSM8K evaluation is conducted under a zero-shot setting. To accommodate the variety of valid answer formats, we extend the strict match patterns provided by the lm-evaluation-harness beyond the standard implementation. Our matching criteria accept both the standard GSM8K format (####) and GSM8K-CoT formats prefixed with “The answer is” or “Answer:”.
@@
Dataset TriviaQA HellaSwag GSM8K GSM-Plus HumanEval HumanEval+ MBPP MBPP+
@@
Code
Reasoning
Code
Reasoning
Language EN EN EN EN EN EN EN EN
# Instances 17,944 10,042 1,319 10,552 164 164 500 378
Few-shot # 4 4 4 (0 for TTC) 5 0 0 3 3
Metric Accuracy Accuracy Accuracy CoT Acc. Pass@1 Pass@1 Pass@1 Pass@1
@@ C.3 GSM8K Overfitting Analysis
Figure 15: Performance versus MoE density after removing GSM8K-related training data. Task loss (top) and accuracy (bottom) are plotted against the ratio of active experts $k$ to total experts $E$ for a fixed active parameter budget. While performance on memorization tasks (TriviaQA, HellaSwag) improves with sparsity, the trend reverses for math reasoning tasks (GSM8K, GSM-Plus) at larger active parameter counts. Dashed segments mark the inverse-scaling regime.
@@ C.5 Detailed Results for Coding Tasks
This appendix provides detailed results for the coding task ablations, mirroring the analyses for mathematical reasoning presented in the main text. Specifically, we detail the non-monotonic relationship between scaling and downstream performance, showing how both task loss (Figure 18, 19) and accuracy (Figure 20) can degrade as pre-training loss improves. We then analyze key architectural factors, including the impact of MoE sparsity (Figure 21) and the optimal Tokens-per-Parameter (TPP) ratio (Figure 23). In addition, we report sparsity–accuracy trends for the coding benchmarks HumanEval+ and MBPP+ (Figure 22).
@@
Figure 21: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, HumanEval and MBPP shift their optima back toward dense models. Task loss against the ratio of active experts $k$ to total experts $E$ for a fixed active parameter budget. In the left two tasks (TriviaQA, HellaSwag), increasing sparsity consistently lowers task loss across all active parameter budgets, in contrast, in the right two tasks (HumanEval, MBPP), once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
Figure 22: At fixed active parameter counts, higher sparsity (lower density) consistently improves performance, but at larger active parameter counts, HumanEval+ and MBPP+ shift their optima back toward dense models. Accuracy against the ratio of active experts $k$ to total experts $E$ for a fixed active parameter budget. Once active parameter counts become large, this trend reverses and denser models begin to outperform their sparser counterparts. Dashed segments mark the inverse‑scaling regime that starts at the black circle; solid segments show the standard scaling region to the right.
@@ C.7 Influence of Optimization Hyperparameter
C.8 Ablation on token budget
To validate our choice of the 125B token budget used in the main experiments, we conducted a preliminary ablation study scaling the training duration to 1T tokens. We compared the baseline model trained on 125B tokens against a model trained on 1T tokens. The training corpus for the 1T run was constructed by upsampling the 973.4B token dataset (Table 1) via random sampling to reach the 1T target. For this experiment, we used a model configuration with width $d=2048$, 16 experts ($E=16$), and top-2 routing ($k=2$). Hyperparameters followed the settings described in Section 3.1, with the exception of the learning rate schedule, where the cosine decay was stretched to accommodate the 1T token duration.
The results are presented in Table 4. Since the extended 1T corpus is predominantly composed of general web text, which serves as a rich source of world knowledge, memorization-intensive tasks such as TriviaQA and HellaSwag exhibited significant performance gains. Conversely, reasoning tasks (GSM8K and GSM-Plus) showed comparatively marginal improvements. We attribute this discrepancy to the scarcity of high-quality mathematics and reasoning corpora available in the open source at the time of experimentation; simply scaling up the volume of general web text contributes minimally to reasoning capabilities. Therefore, extending the training to 1T tokens significantly increases computational costs without yielding proportional gains in reasoning performance, justifying our adoption of the 125B token budget to prioritize examining architectural trade-offs.
Table 4: Performance comparison between models trained on 125B tokens versus 1T tokens. The model configuration is fixed at $d=2048$, $E=16$, and $k=2$.
Token Budget TriviaQA HellaSwag GSM8K GSM-Plus
125B 0.279 0.426 0.318 0.188
1T 0.535 0.538 0.363 0.214
Appendix D The Use of Large Language Models
Use of Large Language Models (LLMs). In preparing this paper, we used large language models only for minor editing tasks, such as improving grammar and clarity of exposition. LLMs did not contribute to the research ideas, methodology, experiments, or analysis, and thus played no substantive role in the scientific contributions of this work.